Tohban Report 2015-05-06

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 29 Apr 2015
End Date: 06 May 2015
Tohban: Nicole Duncan
Tohban email:
Next Tohban:
List all reports


Solar Activity

Most of this week was relatively quiet. Activity picked up yesterday with a series of M-class flares followed by an X2.7 at 22:05 UT. The X flare is not associated with a labeled active region, but can be associated with a magnetically complex region that is just rotating over the NE limb. The complexity of this region is not known at this time, but its likely that activity will pick-up as it rotates on disk.

*We're presently in Normal/Vigorous and I asked for a change to Active/Vigorous following yesterday's event, but this change has not been made yet. 

Note: The hsi_tohban_flare counter routine does not show the X flare in the below table.

How many GOES flares occurred?

 Flares above B, C, M, X class were      12   18     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      8    12     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      7     5     0     0

There were RHESSI flares/GOES flares 214 / 30 over the time range 28-Apr-15 05-May-15

Memory Management

SSR was emptying at the beginning of the week. After yesterday's X flare it did not empty and is now at 12.5%.

Spacecraft Status

Spacecraft operated nominally this week.

Data Gaps


Cold Tip Temperature 1 = 109K 
Cold Tip Temperature 2 = 107K 
Cold Plate Temperature 1 = 131K 
Cold Plate Temperature 2 = 129K 

Temperatures rose this week, as expected.

Detector issues

D2 Front: The light leak issue has gotten worse over the past three weeks for this detector and its effects can be seen in its livetime and fast counts. No action is recommended at this time. Next Tohban should keep an eye on it.

D8 Front: Livetime has been hovering around 95% for three weeks. Its likely a result of fast counts which are around 5,000. No action is recommended at this time.

D8 Rear: Last week, we raised the threshold of D8 rear and this alleviated its elevated fast counts and degraded livetime. It was well behaved for a week but then began to deteriorate once again. Over the course of 3 days its fast counts climbed to 100,000 and livetime dropped to 20%. A threshold change brought it to 10,000 on the fast counts and a livetime around 85%. I asked for another threshold increase, but it hasn't been commanded as of writing this report. This detector is expected to continue to be problematic this week and after speaking to Albert, its recommended that the next Tohban try lowering the voltage.

*Try lowering the voltage on this detector (its currently ~ 4000 V)

D9 Front: Livetime is down to 80%, gradually dropping down from 85% over the course of the week. For the previous two weeks it had stably been around 85%. The fast rates are around 10,000, a rise of 3,000 from the beginning of the week.

*Please raise the front segment threshold from 0x30 --> 0x40. 

Other notes

None to report

Spacecraft Management

Decimation Normal/Vigorous
Night time data (fronts) No nighttime events
Night time data (rears) Taking nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools